OpenStudy (anonymous):

A rectangle has vertices whose coordinates are A(1, 4), B(4, -1), C(-1, -4), and D(-4, 1). Find the coordinates of the midpoints of each side. (Enter answers as fractions.) Midpoint of AB = ( , ) Midpoint of BC = ( , ) Midpoint of CD = ( , ) Midpoint of AD = ( , )

OpenStudy (anonymous):

AB=(5/2,3/2) BC=(3/2,-5/2) CD=(-5/2,-3/2) AD=(-3/2,5/2)

OpenStudy (anonymous):

if A=(x1,y1) and B=(x2,y2) then midpoint of AB = ((x1+x2)/2,(y1+y2)/2)

OpenStudy (anonymous):

man you quys are qood